Author Topic: Science Experiments in the ISS  (Read 94916 times)

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #150 on: 05/10/2009 12:50 PM »
Very interesting outreach article concerning the famous Salmonella experiment, with the increased virulence results obtained, and a possible explanation:

http://science.nasa.gov/headlines/y2009/06may_salmonella.htm

Nickerson's team looked at Salmonella from two shuttle flights to the International Space Station: STS-115 in Sept. 2006 and STS-123 in March 2008. They found that 167 genes were either more or less active in these keyed-up bacteria than in the bacteria that hadn't flown. The team also identified a "master switch" that regulates about one-third of these genes, a protein called Hfq. Activity of this protein was also affected by the low fluid shear conditions of spaceflight.
« Last Edit: 05/10/2009 12:50 PM by eeergo »
-DaviD-

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #151 on: 05/11/2009 04:27 PM »
While Atlantis is getting ready for launch in a great astronomical mission, some ISS updates regarding its science experiments for this past week.

Let's start with a great, long update of BCAT-3/4, which appears to be yielding very interesting observations:

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): Feedback from Peter Lu (PI): Dr. Mike, you did a fabulous job over the weekend setting up and photographing sample 4. The focus and illumination are fantastic, as is clear from the visible phase separation already ongoing (which is, incidentally, much faster than that observed for sample 2 over the past few weeks). And, you managed to trap a large bubble beneath the stir bar, as per the procedures, so it's not interfering with our measurement. Feedback from Chaikin and Hollingsworth at NYU regarding the BCAT-4 crystal growth samples: We are seeing evidence of crystals in samples 8 and 9. Sample 8 is showing some red speckle and curvature in the middle of the sample cell and again to the left. Sample 9 has a small red and green scintillating pattern in the bottom right hand corner. Very nice! Yet, the November pictures of sample 8 and sample 9 looked more dramatic, which indicates that we may want to try taking the photographs at the November angles, since the effects of crystals are likely to be even more pronounced now. Seeing crystals growing in samples 8 and 9 is very exciting and has profound implications regarding how it is that order arises out of disorder in nature. We would be interested to know if you see something similar in sample 10, which has not been captured by the camera. These three samples contain hard spheres, which are monodisperse, polydisperse, and monodisperse seeded with ‘nanodirt’ (large seed particles).”

BISE was also performed as voluntary science this week (there are images of the hardware for this experiment in the previous thread page)

Wakata continues to be very active, performing some slipped JAXA experiments from Expedition 18 (some commercial, some educational), as well as his regularly scheduled activities, like 3DSpace or FACET.

Update on EuTEF's status:

EuTEF (European Technology Exposure Facility):  Platform nominal. DOSTEL, EXPOSE, FIPEX, On-going science acquisition – nominal. MEDET: Acquired science nominally until 5/6 and then was switched off because the instrument was getting too cold. On-going science acquisition - nominal;DEBIE-2: the instrument continues to generate empty science packets at regular intervals (of 30 to 34hrs). Science acquisition is pursued with regular power cycling of the instrument (work-around).

A pair of anomalies in the Protein Crystallization facility in the EDR:

EDR (European Drawer Rack): The rack is continuously active in support of the PCDF (Protein Crystallization Diagnostic Facility) experiment. Two anomalies are currently under investigation.
PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): Growth cycle, EP4 Cycle4 rerun ended on 5/5. In this reactor, science team sees about 10 large crystals that will be kept to be downloaded with PCDF-PU. Additionally in this reactor, the science team observed nucleation occurring at a higher temperature than in ground testing. The cause for this is still being studied.Nucleation cycle EP1 Cycle3 was started on 5/5. This cycle will conclude on 5/8. This cycle studies nucleation at a certain precipitant and protein concentration but at three different temperature rampings.

FSL checkouts continue, with some tests carried out when the last Progress undocked, to check its vibrational modes.
SOLAR is waiting for a Sun observation window to start, and SPICE has been restarted at the end of April.

Finally, a CEO report, and a note concerning difficulties for CEO targets during the next few days or weeks, due to the Station's orbit being almost parallel to the terminator. Lightning conditions are quite bad for Earth Observation purposes.

CEO (Crew Earth Observations): “Through 5/4, the ground has received a total of 5,499 of ISS CEO imagery for review and cataloguing. “Many of your imagery are under review and we are now pleased to confirm your acquisition of views of the following targets: Houston, TX; Delhi, India; Paris, France; Lake Poopo, Bolivia; Lake Chad, Chad; Galapagos Islands, Ecuador; and Kilauea and Mauna Loa, Hawai’i. Nice going! Also you are being very responsive to our specific requests with better than 20% of your imagery being acquired at target times provided. Your excellent, well-focused view of Ankara, Turkey will be posted on NASA/GSFC’s Earth Observatory website this weekend. Among the best recent images from the ISS of the Turkish capital city, your imagery beautifully documents this ancient and complex urban area located in an active earthquake region. Kudos to the crew for a fine image of a requested target!”CEO Note: In recent days, ISS daylight-awake orbit tracks have shifted rapidly into the Southern Hemisphere which is now some six weeks into the fall season, and both day length and sun elevation are significantly lowering. This situation along with deteriorating seasonal weather greatly limits good view opportunities for targets. The situation is further aggravated by the fact that the ISS orbit tracks nearly parallel with the terminator. The consequence is very low light right of track, low light near nadir, and adequate to good light left of track. Beginning today and for the next 5 to 7 days, there may be no targets with suitable illumination or weather.[/i][/i][/i][/i][/i][/i]
« Last Edit: 05/11/2009 04:29 PM by eeergo »
-DaviD-

Offline Marsbug

  • Member
  • Posts: 43
  • Liked: 0
Re: Science Experiments in the ISS
« Reply #152 on: 05/12/2009 02:28 PM »
Hell, could some one please help, I was wondering if results from the previous BCAT experiments been published in any journals? I've been looking on the NASA site but I keep getting 404 errors.
'We came in peace, for all Mankind' Consider that phrase and all it implies.

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #153 on: 05/12/2009 10:13 PM »
Hell, could some one please help, I was wondering if results from the previous BCAT experiments been published in any journals? I've been looking on the NASA site but I keep getting 404 errors.

Not a science journal, but this outreach explanation of the experiment is quite helpful to get a clear idea of supercriticality and its applications:

http://www.nasa.gov/mission_pages/station/science/BCAT_feature_093005.html

For more, in-depth information about this experiment, check the updated GRC site:

http://spaceflightsystems.grc.nasa.gov/Advanced/ISSResearch/MWA/BCAT/

More to the point of what you're looking for (from the NASA site, perhaps you disregarded them):

http://www.opticsinfobase.org/abstract.cfm?URI=AO-40-24-4146
http://www.ph.ed.ac.uk/~abs/PhysRevLett_96_028306_2006.pdf
http://scitation.aip.org/getabs/servlet/GetabsServlet?prog=normal&id=APCPCS000458000001000108000001&idtype=cvips&gifs=yes

If you're interested in more but don't know how to find them, I'm just googling the references in NASA's site http://www.nasa.gov/mission_pages/station/science/experiments/BCAT-3-4-CP.html and looking for a relevant journal that contains the abstract. Of course, to get the full text you'll have to pay for it, or get the journal in a library.

If you still need more details but don't want to get into full-blown science papers, consider writing an email to the experimenters. See the GRC site, in the front page, at the bottom, to find their respective addresses.
-DaviD-

Offline Marsbug

  • Member
  • Posts: 43
  • Liked: 0
Re: Science Experiments in the ISS
« Reply #154 on: 05/13/2009 09:46 AM »
Thank you eeergo, that's very helpfull! :)
« Last Edit: 05/13/2009 09:46 AM by Marsbug »
'We came in peace, for all Mankind' Consider that phrase and all it implies.

Offline chen2x18

  • Member
  • Posts: 1
  • Liked: 0
Re: Science Experiments in the ISS
« Reply #155 on: 05/18/2009 04:25 AM »
It is hard core engineering and science and a functional system because of what experiments they gathered..

_________________
Micro molding



Offline robertross

  • Canadian Member
  • Senior Member
  • *****
  • Posts: 16338
  • Liked: 54
  • Westphal, Nova Scotia
Re: Science Experiments in the ISS
« Reply #156 on: 05/24/2009 12:02 AM »
Okay, my turn!

ISS On-Orbit Status 05/23/09

All ISS systems continue to function nominally, except those noted previously or below. Saturday – light-duty day for the ISS crew (soon to double in size!).

FE-1 Barratt continued his second run of sleep logging for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. This is similar to Barratt’s BCD (Baseline Data Collection) which was performed pre-flight for comparison. [To monitor the crewmember’s sleep/wake patterns and light exposure, Mike wears a special Actiwatch device which measures the light levels encountered by them as well as his patterns of sleep and activity throughout the Expedition and uses the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

The crew performed the regular weekly three-hour task of thorough station cleaning, including COL (Columbus Orbital Laboratory) and Kibo. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, CDR Padalka conducted regular maintenance inspection & cleaning on fan screens in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9), plus dust filter replacement in the FGB.

The CDR also performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

For his chosen VolSci (Voluntary Weekend Science) program, FE-2 Wakata worked in the Kibo JPM (JEM Pressurized Module), installing dehumidifier units in the CBEF (Cell Biology Experiment Facility) incubator’s micro-G & 1-G sections.

FE-1 Barratt’s VolSci choice, troubleshooting the nonfunctional AgCam (Agricultural Camera), was to look into two failure possibilities identified on the ground team’s fault tree: (1) Bent or broken pins on the AgCam data acquisition cable (straighten pins if bent, wait for a new cable if broken), or (2) failed parallel port on the AgCam laptop (if so, swap with ELT-3 {EXPRESS-3 Rack laptop}). Further steps depend on the outcome of the troubleshooting.

In the US Airlock, Wakata wrapped up the BC4 (Battery Charger 4) troubleshooting by terminating the test discharge of EMU (Extravehicular Mobility Unit) battery #2037 in BC4.

Koichi also completed the regular bi-monthly reboots of the OCA Router and File Server SSC (Station Support Computer) laptops.

At ~9:00am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

At ~11:55am, Gennady Padalka powered up the SM's amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and at ~12:00pm conducted a ham radio session with the participants of the Second Russian Innovation Forum at Kursk.

The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR), ARED advanced resistive exercise device (CDR, FE-1, FE-2) and VELO cycle ergometer with bungee cord load trainer (FE-1, FE-2). [The CEVIS (Cycle Ergometer with Vibration Isolation) continues to be unusable (see above).]

Afterwards, Mike downloaded the exercise data file to the MEC (Medical Equipment Computer) for downlinking, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

CDR Padalka had four job items on his discretionary “time permitting” task list:

A session for Russia's Environmental Safety Agency (EKON), making observations and taking KPT-3 aerial photography of environmental conditions on earth using the Nikon D2X with the SIGMA 300-800mm telephoto lens.
Another run with the GFI-8 "Uragan" (hurricane) earth-imaging program, using the NIKON D2X digital camera to take 800mm-lens telephotos for subsequent downlinking on the BSR-TM payload data channel,
Urine transfer to the Progress Rodnik tankage, and
An audit/inventory of TP-TRG-L thermally conducting gaskets for various electronic components such as GIVUS A6, the TVM & TsVM Computers, and SNT Voltage & Current Stabilizers.

Conjunction Advisory: No maneuver required.

STS-125/Atlantis Return: Due to bad weather, today’s return of Atlantis to KSC was called off for all opportunities.
Return opportunities tomorrow (Sunday) are (EDT):
· 10:11am (Orbit 196) at KSC (Deorbit Burn: 8:58am)
· 11:40am (Orbit 197) at EDW (DO: 10:25am)
· 11:49am (Orbit 197) at KSC (DO: 10:31am)
· 1:19pm (Orbit 198) at EDW (DO: 12:08pm)

Weekly Science Update (Expedition Nineteen -- Week 7)

3-D SPACE: Ongoing.

AgCam (Agricultural Camera): Planned.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): “Nice focus. Very interesting images of sample 5, phase separation sample, which merit additional thought. Promised comment last week on crystal samples. Sample 8 (15830) showing crystals, but not as colorful as expected. May need to reduce reflections to see vibrant colors; still pondering and grateful as always for your nice work. Thank you!”

BIOLAB (ESA): No report.

Biological Rhythms (JAXA): Ongoing. Two runs are for Barratt, two are planned for De Winne.

BISE (CSA, Bodies in the Space Environment): Ongoing.

BISPHOSPHONATES: “Koichi, thank you for continuing to take your Alendronate pill. Your next dose will be Monday (5/25). Please continue to adhere to the fasting and NSAID constraints. Thanks for supporting this research.”

CARD (Long Term Microgravity: A Model for Investigating Mechanisms of Heart Disease, ESA): “Many, many thanks Koichi for your dedication and efforts to perform this experiment on 5/20 and 5/21. The CARD protocol includes a 24hrs urine collection, a 24hrs blood pressure monitoring, a blood draw (in the morning of day#2), and 5 cardiac output measurements performed with the HRF-2 Pulmonary Function System (PFS) via re-breathing technique. On 5/20, most of the CARD activities have been performed. An anomaly with the CARDIOLAB blood pressure Holter was encountered and the instrument could not be programmed for the blood pressure monitoring overnight. On 5/21, blood draw, urine collection and two PFS sessions have been completed nominally. Unfortunately, the Holter anomaly forced ESA to call off the continuation of the blood pressure monitoring in conjunction with the remaining two PFS sessions on 5/21. Moderate science impact is expected. Ground team will try to troubleshoot the Holter ASAP. ESA could envisage the possibility to plan for the blood pressure monitoring overnight later on if the Holter anomaly gets solved.”

CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Mike, your next CCISS session is planned for Wednesday (5/27) and Thursday (5/28).

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): “CIR ran 3 FLEX test points on Friday, 5/22. These were the first of the 196 FLEX test points that are planned. We will not know about the success of those test points until after we are able to downlink the images, which isn’t on the timeline until next week, tentatively scheduled for 5/28.”

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Cosmo Flower mission was completed by Wakata (slipped from I-18)

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the Protein Crystallization Diagnostic Facility (PCDF) experiment. Over the last couple of weeks, some power trips and reboots of EDR and/or PCDF have been encountered, and engineering teams are pin-pointing a flimsy EDR ESEM Board as the root cause of these problems. ESA plans to perform the swap of this ESEM Board during Week#8 with a spare available on orbit. To perform the swap, EDR and PCDF will have to be powered down for 1h30m minimum (impact to science to be minimized with respect to planned runs and previously formed crystals).

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

ENose (Electronic Nose): Operating.

EPM (European Physiology Module): Underway.

EPO (Educational Payload Operations): Reserve.

EPO J-Astro Report (JAXA): “According to EPO request, total report number has increased twofold. Public readers expecting your new reports. We appreciate interesting Space report; please continue with your outstanding job.”

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): Complete.

EPO Moon Score (JAXA): Planned.

EPO Try Zero-G (JAXA): Planned.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): “Additional photo was very interesting. Thank you. Please keep photo in proprietary folder when you download them.”

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): EuTEF platform and its instruments have been safe-configured for the NH3 venting on 5/15. - DEBIE-2: The last EOP script has been nominally started on 5/17 and will run for ~9 days. The instrument continues to regularly generate empty science packets, and as a work-around the instrument script includes a daily power cycle command, which is skipped from ground if the science packets remain nominal. - DOSTEL: On-going science acquisition - nominal;- EXPOSE: On-going science acquisition. On 5/13, the valves of the Experimental Trays have been closed for the NH3 venting event. It is planned to keep those closed for about 2 weeks to avoid any contamination to the samples;- EVC: Instrument is too cold to be operated for the time being;- FIPEX: The last EOP script has been nominally started on 5/17 and will run for ~9 days; - MEDET: the instrument has been manually restarted on 5/19;- PLEGPAY: Inactive, “Experiment 1” memory has been erased on 10/30/2008. Plasma generation capability has been disabled. On 5/21, PLEGPAY Langmuir Probe measurements have been resumed (after ~8.5mths of stand-down due to safety concerns).

FACET (JAXA): “Facet experiments continue smoothly. Microgravity environment in JEM seems very good for the experiment.”

FSL (Fluid Science Laboratory): No report.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): “#2 for FE-2 was completed. There may be low signal intensity for Ch2, but data analysis will be no problem. We believe this is good first step for next Biorhythms experiment as well as future telemedicine demonstration.”

HQPC (JAXA): To be launched by 34P.

ICE CRYSTAL (JAXA): Complete.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

Integrated Immune: Complete.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

MAXI (JAXA): “This is not utilization but really close to (External) Payload activity, thus included in this week’s report: .Step 1 is complete and we confirm that modified MAC address data can correctly reach SSIPC. Good step for MAXI to recover from MAC address issue and start operations after 2JA. We need your contribution for Step 2 and 3 before 2JA.”

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: Ongoing.

PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): Over the last weekend, the last nucleation runs have been successfully performed with EP1 reactor. EP1 Cycle6 has been successfully performed from 5/13 to 5/16. At the end of this run, a minor anomaly has occurred with the PCDF Light Scattering Unit (LSU) device, resulting in a loss of the last 2hrs of Dynamic Light Scattering measurements. Since 5/16, science team is now processing EP3 reactor, with similar nucleation runs, but for a different range of protein concentration. EP3 Cycle1 has been started on 5/16 but had to be aborted due to a PCDF spontaneous reboot on 5/17. EP3 Cycle1 rerun has been restarted and performed successfully from 5/17 to 5/19. EP3 is running nominally since 5/19. The EP4 reactor which contains some nice crystals to be returned by 2J/A. In preparation for the ESEM Board exchange, the science team has decided to progressively decrease its temperature from +23degC to +20degC by 1degC steps (over 24hrs) with some days in-between the temperature decreases. Finally, several important aspects of the remaining PCDF mission objectives have been discussed this week with the science team: 1) reshuffling of remaining science runs until 2J/A return; 2) coordination of periods of interest for SAMS measurements; 3) contingency plan in case of problem with EDR ESEM Board exchange.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

POLCA/GRAVIGEN (ESA): Complete.

RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Koichi, next week we are scheduling your 2nd week of Sleep logging for this increment. Mike, thanks for completing the monthly Actiwatch download/initialization session and getting the watches ready for Frank and Bob. The watches will take data until 6/14.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): The next Sun visibility window is expected to start on 5/23. The platform has been put a safe mode for the 33P docking on 5/12. For the NH3 venting event on 5/15, the platform has been put in idle mode (i.e. motor-controlled and fixed) and all the instrument covers closed. After the NH3 venting, SOLAR was put in survival power (feeder#2 only) configuration.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): Planned.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

ISS Orbit (as of this morning, 8:32am EDT [= epoch])
Mean altitude -- 350.3 km
Apogee height – 356.9 km
Perigee height -- 343.7 km
Period -- 91.54 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0009804
Solar Beta Angle -- 24.7 deg (magnitude peaking)
Orbits per 24-hr. day -- 15.73
Mean altitude loss in the last 24 hours -- 76 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 60206
A new season of Doctor Who is fast approaching!!

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #157 on: 05/29/2009 11:36 AM »
Thanks for the update Robert! :)

Science@NASA has just published a new article on Matroshka's results and its implications for long duration spaceflight, especially outside the magnetosphere. Surprisingly, even though the current models are ok within 10%, NASA doesn't appear to have a sufficiently effective radiation shielding system to support Mars missions, and would barely stay within safety limits in 6-month lunar missions. It appears not everything's about rockets in going 'beyond' ;)

http://science.nasa.gov/headlines/y2009/27may_phantomtorso.htm
-DaviD-

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #158 on: 06/02/2009 09:40 PM »
This week's update... these are hopefully going to get much bigger with the brand new 6-person crew. Some of the experiments that are to begin soon thanks to the newly arrived deWinne, Thirsk or Romanenko are Biological Rythms from JAXA, BISE from the CSA, and WAICO from ESA.

First of all, some anomalies noted in AgCam, but nothing serious:

AgCam (Agricultural Camera): “Mike, the AgCam team was very happy with your successful laptop exchange during your Voluntary Science activity on Saturday, and were all geared up to get good imagery this week. However on Wednesday, 5/27, when attempting to take their first image, another anomaly, unrelated to the first one, became apparent. The AgCam team believes a short troubleshooting crew activity to change a cable would do the trick, and a task list activity to do that is under development.”

JAXA's cell biology experiment is having some difficulties getting set up, but is planned during the 17A stage, so there's still time.

CBEF (JAXA Cell Biology Experiment Facility): Troubleshooting continues. “The CBEF dehumidification was completed. Humidity reached to certain level to be able to start data acquisition of the CO2 sensor which has sensitive for humidity. Fan Rotations have also been finished by Koichi. Thank you for providing more effort than the scheduled task. The results show that the fan of the temp control unit in the micro gravity compartment is completely failed. The ground engineering team is now investigating how to repair this function. The first repair task will be prepared and need to be performed before running the Space Seed experiment, currently scheduled during the 17A stage.”

There have been also some problems on the heptane combustion runs in the CIR, and in a single propane run:

CIR (Combustion Integrated Rack): During last week’s six test point attempts (three methanol, three heptane), two methanol burns were successful but only one of the heptane burn attempts was partially successful. The first methane droplet drifted into the igniter, which immediately stopped the burn. The first two heptane attempts did not burn while the third heptane droplet hung on the dispenser needle and did not have a true microgravity characteristic during its burn. The heptane test point will be re-run at TBD date. The team is still assessing future operations with the given stowage configuration around the CIR rack.

FACET is doing well and will soon be completed:

FACET (JAXA): “75% of experiments have finished. Data show that Microgravity condition in JEM is really good for FACET experiment. We planned two more weeks to run the experiment and all will finish just before the 2J/A arrival if everything goes well.”

The Microgravity Measurement Apparatus in the JEM has completed its first run, and is awaiting Endeavour for data retrieval and the start of its second run:

MMA (Microgravity Measurement Apparatus): First attempt using ELT (Experiment Laptop Terminal) with MMA for micro-G data acquisition was successfully finished after a laptop exchange from MLT (MMA Laptop Terminal) to ELT. We successfully acquired background microgravity data in the JEM using the modified configuration. The MLT hard drive was removed by Koichi (thank you for quick task) and expected to be retrieved by 2J/A for micro-G data salvage on the ground. New hard drive installation will be scheduled after 2J/A which will carry a new drive unit.

Some good news from EuTEF, which is at last operating at its full potential. Specifically, while EXPOSE did temporarly close its lids to avoid ammonia contamination from the S3 radiator venting, PLEGPLAY has reactivated its Langmuir probe, the one that caused concerns about plasma discharge and the Soyuz pyrobolts. I am guessing the Pirs Langmuir probe gives enough confidence now for its activity to resume. I'm not really sure about the status of TRIBOLAB, which was experiencing some difficulties the last time I checked.

EXPOSE: On-going science acquisition. On 5/13, the valves of the Experimental Trays have been closed for the NH3 venting event. It is planned to keep those closed for about 2 weeks to avoid any contamination to the samples;-- PLEGPAY: On 5/21, PLEGPAY Langmuir Probe measurements have been resumed (after ~8.5mths of stand-down due to safety concerns). Preliminary analysis of the measurements shows that the Langmuir Probe is in good shape. Science measurements have been resumed between 5/26 and 5/28. PLEGPAY instrument will be de-activated one day prior to 19S docking.

Finally, very complete update about EDR and its resident experiment, PCDF. In spite of some temperature problems, it appears science impact will be restriced, and some samples are to be returned during STS-127:

EDR (European Drawer Rack, ESA): “Thank you Mike for having successfully performed the exchange of the ESEM Board. During this activity, the PCDF instrument was powered down for 2hrs in total, which is short enough from a science point of view. In order to minimize the temperature fluctuations for the PCDF reactors during the planned power down period, Columbus Moderate Temperature Loop (MTL) set-point was increased from +17degC to +19degC. While the ESEM showed good H&S data from the re-activation of EDR, some anomalies have been encountered during PCDF restart (see dedicated PCDF paragraph). The EDR status is fully nominal since 5/27.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): Over the last weekend, one additional science run has been performed with the EP3 reactor from 5/22 to 5/24. On 5/25, science acquisition started with the last of the 4 PCDF reactors, i.e. EP2. This last reactor is dedicated to study the phenomena of protein depletion zone apparition during crystal growth phase. Initial measurements had to be interrupted in order to perform the EDR ESEM board exchange on 5/27. Both PCDF Electronics Unit (EU) and PCDF Process Unit (PU) were powered off for a total duration of 2hrs during this EDR ESEM board exchange. After this crew activity, EDR got re-activated, followed by PCDF-EU and PCDF-PU power on sequence. Very rapidly an anomaly occurred during the PCDF-PU boot-up, as it started to rapidly cool down in an unexpected way (science impact TBC as soon as new images are available on ground). Ground team immediately performed a power cycle of the PCDF-PU, and normal temperature control of the PCDF-PU was then recovered, but leading to different operative mode of the PU (i.e. Stand-Alone Mode), which did not allow to resume the science acquisition. PCDF was kept in this configuration overnight, with a higher temperature set-point of the Columbus Moderate Temperature Loop (MTL) at +19degC (instead of usual +17degC). On 5/28, ESA engineering experts gave their recommendation to proceed with a full and clean power cycling of both PCDF-EU and PCDF-PU. After ground commanding, PCDF-PU was put back in a nominal configuration and the science acquisition could be restarted. The EP4 reactor still contains some nice crystals to be returned by 2J/A. Next weeks will be devoted to grow additional crystals in EP2 reactor.

Oh boy, is this text formatter driving me crazy... sorry for the misplaced italics or strange symbols, I keep editing the post format and it keeps getting all screwed up.[/i][/i]
« Last Edit: 06/02/2009 09:46 PM by eeergo »
-DaviD-

Offline robertross

  • Canadian Member
  • Senior Member
  • *****
  • Posts: 16338
  • Liked: 54
  • Westphal, Nova Scotia
Re: Science Experiments in the ISS
« Reply #159 on: 06/03/2009 01:49 AM »
Thanks David!

Wow, they finally have EuTEF going again!!!!!! Awesome news. Quite the troublesome unit...although for ISS that's nothing new.

And stowage issues around the CIR. They really need another module up there to move some stuff around (preferrably permanent MPLM).
A new season of Doctor Who is fast approaching!!

Offline erioladastra

  • Full Member
  • ****
  • Posts: 1115
  • Liked: 15
Re: Science Experiments in the ISS
« Reply #160 on: 06/04/2009 01:54 AM »
Thanks David!

Wow, they finally have EuTEF going again!!!!!! Awesome news. Quite the troublesome unit...although for ISS that's nothing new.


I think you might be confused.  EuTEF has been working fine for some time.  One of many experiments, the Langmuir probe, was deactivated for awhile over safety concerns.  But the other experiements are doing well.

Offline catfry

  • Member
  • Full Member
  • ****
  • Posts: 402
  • Liked: 0
Re: Science Experiments in the ISS
« Reply #161 on: 06/04/2009 05:27 AM »
There was also one of the payloads that tended to send empty data packets.

Offline robertross

  • Canadian Member
  • Senior Member
  • *****
  • Posts: 16338
  • Liked: 54
  • Westphal, Nova Scotia
Re: Science Experiments in the ISS
« Reply #162 on: 06/05/2009 08:56 PM »
Thanks David!

Wow, they finally have EuTEF going again!!!!!! Awesome news. Quite the troublesome unit...although for ISS that's nothing new.


I think you might be confused.  EuTEF has been working fine for some time.  One of many experiments, the Langmuir probe, was deactivated for awhile over safety concerns.  But the other experiements are doing well.

I must have missed that update then...thanks.
A new season of Doctor Who is fast approaching!!

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #163 on: 06/07/2009 03:43 PM »
This week's update:

BCAT is proving to be very popular among astronauts, with many extras being performed. The science team is also very "talkative", at least in public releases:

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): “Thank you Dr. Mike for the bonus pictures of sample 5. It’s really great to get this extra science, along with your advice. We are working on implementing your suggestion to move BCAT to the end-cone of node-2. And thanks also for offering to take more photographs of the crystal samples 8, 9, 10. When we saw the big colourful crystals in the early CDOT experiment many years ago we were using cylindrical sample cells. Since we can only form these entropy-driven crystals in microgravity (i.e. a more ordered structure gives the particles more room in which to rattle about, or higher entropy), it was important to learn from you that the better colour separation that should result from using rectangular sample cells is challenging to capture with a camera. Thanks again.”

As those of you who watched the pre-EVA22 press conference may remember, some the Exp20 crewmembers who were in the US segment during the spacewalk (DeWinne, Romanenko and Thirsk) performed some BISE runs. Apart from that, this week's update:

BISE (CSA, Bodies in the Space Environment): “Thanks Bob for your excellent work during your first BISE session. The data has been validated and the PI team and CSA staff are excited by the results. Thanks Frank for excellent work too and thank you both for a very creative way to reposition the laptop and the camera. Your data has been validated and is also interesting.”

Some CBEF news, spotlightning the CO2 sensor calibration.

CBEF (JAXA Cell Biology Experiment Facility): Troubleshooting continues. “CO2 sensor data were measured under the following condition: 37degreeC, RH 35%, 40%, 45%. Next step will be to dry inside the CBEF by desiccant and measure RH20%-35%. Then we can make a calibration curve for the CO2 sensor.”

There are many JAXA educational experiments planned, completed or in reserve for Koichi, who appears to be very active in this regard (well, and in many others, he always appears several times in the news reports, and TVIS has been keeping him busy for some weeks now):

EPO (Educational Payload Operations): Reserve.
EPO J-Astro Report (JAXA): Ongoing.EPO Space Clothes (JAXA): Complete.EPO Hiten (Dance, JAXA): Complete.EPO Moon Score (JAXA): Planned.EPO Try Zero-G (JAXA): “#3 (Twosome) will be Voluntary Science on this weekend.”EPO Kibo Kids Tour (JAXA): Complete.EPO Spiral Top (JAXA): “Additional photo was very interesting. Thank you. Please keep photo in proprietary folder when you download them.”

There are also some EuTEF news to expand on the last few posts. The shutters in EXPOSE are being kept closed for 2 weeks from the date of the ammonia venting. We also have an update on TRIBOLAB's 'mystery' status:

[
i]T
RIB
OLA
B:
was
co
mmand
ed into thermal stabilization mode on 6/4 and an integrity check of the motor is planned for 6/5.
 
[/i]ICV (cardiovascular experiment) is about to start this coming week:

ICV (Integrated Cardiovascular
):
The
IC
V t
eam
is
look
ing forward to starting on-orbit ops next week!

The JPM microacceleration measurement apparatus (MMA) is being checked by Koichi, by tapping with is fist on the JPM's windows, and seeing how the vibrations affect the instrument's readings:

MMA (Microgravity Measurement Apparatus):  “Micro-g measureme
nt
was
pe
rfo
rme
d t
o inv
estigate how microgravity disturbance will transmit to the Experiment Rack in JEM. Data are now being analyzed. Koichi, we caught your tapping signal by the MMA sensor, and it is good reference for analysis.”

Another experiment that was performed during the spacewalk inside the USOS was NeuroSpat. There are photos of this cognitive experiment in ESA's site (for example http://www.esa.int/esa-mmg/mmg.pl?b=b&keyword=neurospat&single=y&start=3)

NEUROSPAT (ESA): First sessions were performed by Bob & Frank on 6/3 and 6/4 respectively. “We wou
ld
lik
e t
o t
han
k t
he
both of you for your dedication and hard work at getting all the science objectives met. Thank you for all the excellent ideas that made it possible to complete the execution of this experiment. The science team is very much looking forward to receiving the data and starting their analysis.”

Some more protein crystallization fun (there's always some anomaly in this payload!):

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): The nucleation run on EP3 Cycle 4 ended on 5/31 and the second c
ycl
e o
f E
P2
was
st
arted
on the same day. This reactor is dedicated to study the phenomena of protein depletion zone apparition during crystal growth phase. The science team observed crystals and depletion zones in this cycle. MMA measurements were also performed in support of this run. On 6/4 a PCDF anomaly occurred, and PCDF spontaneously rebooted and was off for about 3 minutes and the contingency scripts started. The science team made the decision to re-dissolve and start the cycle again, with an adjustment to the temperature profile. The rerun of the cycle was started 2 hours after the spontaneous reboot. The EP4 reactor still contains some nice crystals to be returned by 2J/A. The next week will be devoted to grow nice crystals in EP2 reactor.

Getting up to date in SLEEP:

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Koichi, thanks for completing another week of sleep logging. We currently have you scheduled to co
mpl
ete
an
oth
er
wee
k nex
t week. This will actually be your fourth total due to a loss of data from the SSC during one of your previous sleep logging weeks. Mike, next week we are scheduling another week of sleep logging. Thanks for your participation. Bob & Frank, you are currently completing your first week of sleep logging. We are targeting every third week for future sleep logging.”

SOLAR's observation window has ended on June 5th. No word on when it'll be reestablished yet.

EDIT: And, one more day, my regularly scheduled problems with the text editor :D Please, disregard poor editing.[/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i][/i]
« Last Edit: 06/07/2009 03:49 PM by eeergo »
-DaviD-

Offline eeergo

  • Phystronaut
  • Senior Member
  • *****
  • Posts: 3857
  • Liked: 35
  • Blacksburg, VA; Italy; Spain
Re: Science Experiments in the ISS
« Reply #164 on: 06/07/2009 03:48 PM »
And also the second part of the Matroshka article I noted here last week online at Science@NASA:

http://science.nasa.gov/headlines/y2009/03jun_fakeastronaut.htm

Now, it focuses more on the 'ground' segment of this experiment, where they will subject a Matroshka model to an artificial 'solar storm' event.
-DaviD-

Tags: